OpenStudy (linn99123):

Which is not a solution to the equation -3x+y=5 A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8) I picked C

OpenStudy (solomonzelman):

plug in your options.

OpenStudy (texaschic101):

-3x + y = 5 (-4,-7) x = -4 and y = -7 -3(-4) + (-7) = 5 12 - 7 = 5 5 = 5 (this is correct, so it is not this one) (3,14) x = 3 and y = 14 -3(3) + 14 = 5 -9 + 14 = 5 5 = 5 (correct, so not the right one) (2,-1) x = 2 and y = -1 -3(2) + (-1) = 5 -6 - 1 = 5 -7 = 5 (BINGO..here it is) (1,8) x = 1 and y = 8 -3(1) + 8 = 5 -3 + 8 = 5 5 = 5 (correct...not this one) you got it right...answer is C